# Courcelle’s Theorem for Lipschitz Continuity

### v1

@@ 7 Clique-Width
In this section, we prove Theorems 1.6 and 1.7. Since the proofs are similar to that for Theorem 1.6, we provide most of the proof of Theorem 1.7 in Appendix C. Let $G$ be a vertex-colored graph with $n$ vertices. A VR-term $\mathcal{T}$ of $G$ represents a procedure that constructs $G$ by recursively applying the following operations [CourcelleE2012]:
@@ 7.1 Definitions
for minimization problems. If $\mathsf{sat}(G,\varphi)=\emptyset$, we define $\mathsf{opt}_{w}[G,\varphi]=\mathtt{nil}$. For graphs $G$ and $\textsf{MSO}_{1}$ formulas $\varphi$ with $\mathsf{sat}(G,\varphi)\neq\emptyset$, our algorithm recursively computes the vertex set $\mathsf{DP}_{w}[G,\varphi]$ that approximately achieves the maximum in Equation (5) or minimum in Equation (6). When it is clear from the context, we omit the subscript $w$ and simply write $\mathsf{opt}[G,\varphi]$ and $\mathsf{DP}[G,\varphi]$. We perform the dynamic programming algorithm using the formulas defined in Theorem 7.2 over the parse tree of a term obtained from Lemma 7.1. Specifically, at every stage of the algorithm, it is guaranteed that each $X\in\mathsf{sat}(G,\varphi)$ satisfies $X\subseteq V(G)$. Furthermore, our algorithm is randomized. Thus, $\mathsf{DP}[G,\varphi]$ can be considered as a probability distribution over vertex sets in $\mathsf{sat}(G,\varphi)$.
@@ 7.4 Putting Together
Here we complete the proof of Theorem 1.6. Let $G=(V,E)$ be a graph with clique-width $\mathrm{cw}$. From Lemma 2.1, we can compute a VR-term $\mathcal{T}$ over a $k$-colored graph denoting $G$ with height $O(\log n)$. We perform the dynamic programming algorithm on the parse tree of $\mathcal{T}$. We remark here that operations other than disjoint union do not affect either the approximability or the Lipschitz continuity, by setting
@@
Proof 7.6.
The approximability bound is obtained by repeatedly applying Lemma 7.3 and $(1-\varepsilon)^{h}\geq 1-h\varepsilon$. The Lipschitzness bound is obtained by repeatedly applying Lemma 7.4.
@@
Proof 7.7 (Proof of Theorem 1.6).
The claim follows by substituting $\varepsilon$, $h$, and $k$ in Lemma 3.9 with $\frac{\varepsilon}{\Theta(\log n)}$, $O(\log n)$, and $O^{*}(2^{k})$, respectively.
@@ Appendix A Omitted Proofs on Logic
It is easy to verify that the resulting parse tree represents $G$, and the height is at most $O(\log k+h)$.
Proof A.2 (Proof of Corollary 2.6).
First we show the first statement of Corollary 2.6. Since $S^{\prime}\subseteq B\subseteq\mathsf{src}(G)$, $\mathsf{sat}(G,\psi\mathord{{\restriction}_{S\cup S^{\prime}}})$ and $\{S^{\prime}\}$ is separated. It is clear that families of vertex sets $(\mathsf{sat}(G,\psi\mathord{{\restriction}_{S\cup S^{\prime}}})\boxtimes\{S^{\prime}\})_{S^{\prime}\subseteq B}$ are disjoint.
Let $\psi(X)$ be a $\textsf{C}_{r}\textsf{MSO}^{q}$-formula obtained by applying the first statement of Theorem 2.5 to $\varphi(X)$ and $B$. Then, $\mathsf{fg}_{B}(G)\models\varphi(A)\iff G\models\psi(A)$ for any vertex set $A$. We show that $\mathsf{sat}(\mathsf{fg}_{B}(G),\varphi\mathord{{\restriction}_{S}})=\biguplus_{S^{\prime}\subseteq B}\mathsf{sat}(G,\psi\mathord{{\restriction}_{S\cup S^{\prime}}})\boxtimes\{S^{\prime}\}.$
@@
Next, we show the second statement of Corollary 2.6. From Theorem 2.5, there exists a family of tuples $\{\theta^{\prime}_{i}({X}),\psi^{\prime}_{i}({X})\}_{i\in[p^{\prime}]}$ of $\textsf{C}_{r}\textsf{MSO}^{q}$-formulas with a free variable ${X}$ such that, for any $k$-graphs $G$ and $H$,
@@ Appendix B Analysis of Softmax and Softmin
$\mathsf{TV}\left(\bm{c},\bm{c}^{\prime}\right)=\Pr\left[\bm{c}\in\left[\frac{2q}{x+\delta},\frac{2q}{x}\right]\right]=\frac{\frac{2q}{x}-\frac{2q}{x+\delta}}{\frac{2q}{x}-\frac{q}{x}}=2-\frac{2x}{x+\delta}=\frac{2\delta}{x+\delta}\leq\frac{2\delta}{x}.$
@@
$=\left(1-\frac{\epsilon}{2}\right)\cdot\left(1-\frac{\epsilon}{2}\right)\max_{i\in[p]}x_{i}\geq(1-\epsilon)\max_{i\in[p]}x_{i}.$
Proof B.4 (Proof of Lemma 2.9).
@@
$\mathsf{TV}\left(\mathsf{softargmax}^{\epsilon}_{i\in[p]}x_{i},\mathsf{softargmax}^{\epsilon}_{i\in[p]}x_{i}^{\prime}\right)\leq\frac{2\delta}{\max_{i\in[p]}x_{i}}+\frac{8\epsilon^{-1}\log(2p\epsilon^{-1})\delta}{\max_{i\in[p]}x_{i}}\leq\frac{10\epsilon^{-1}\log(2p\epsilon^{-1})\delta}{\max_{i\in[p]}x_{i}}.$
@@
$\leq\left(\left(1+\frac{\epsilon}{2}\right)+\left(\frac{\epsilon}{2}\right)^{2}\cdot 2\right)\cdot\min_{i\in[p]}x_{i}\leq(1+\epsilon)\min_{i\in[p]}x_{i}.$
Proof B.6 (Proof of Lemma 2.11).
@@
$\mathsf{TV}\left(\mathsf{softargmin}^{\epsilon}_{i\in[p]}x_{i},\mathsf{softargmin}^{\epsilon}_{i\in[p]}x_{i}^{\prime}\right)\leq\frac{2\delta}{\min_{i\in[p]}x_{i}}+\frac{8\epsilon^{-1}\log(2p\epsilon^{-1})\delta}{\min_{i\in[p]}x_{i}}\leq\frac{10\epsilon^{-1}\log(2p\epsilon^{-1})\delta}{\min_{i\in[p]}x_{i}}.$
@@ C.1 Treewidth
Here, we provide the minimization version of the framework from Section 3. The algorithm is obtained simply by replacing $\mathsf{softargmax}$ with $\mathsf{softargmin}$. Some parts of the analysis are identical to that in Section 3. Specifically, for the introduce operation, exactly the same results can be obtained with exactly the same proof. Furthermore, by replacing $(1+\epsilon)$ with $(1-\epsilon)$ and reversing all inequalities in the proofs of Lemmas 3.1 and 3.5, we obtain the following.
Lemma C.1.
@@
$\leq 30\alpha\epsilon^{-1}\log(2p\epsilon^{-1})\delta+\beta+\delta\leq 31\alpha\epsilon^{-1}\log(2p\epsilon^{-1})\delta+\beta.$
The following lemma corresponds to Lemma 3.9, but slightly different.
Lemma C.7.
@@
Proof C.8.
The approximability bound is obtained by repeatedly applying Lemmas C.1 and C.2. The Lipschitzness bound is obtained by repeatedly applying Lemmas C.3 and C.5 together with $\alpha\leq(1+\epsilon)^{h}$.
@@
Proof C.9 (Proof of Theorem 1.2).
We prove the theorem by substituting $\epsilon$, $h$, and $t$ in Lemma C.7 with $\frac{\epsilon}{O(\log n)}$, $O(\log n)$, and $23t-$ 2, respectively. Let $c\in\mathbb{R}_{>0}$ be the number such that $c\log n$ is an upper bound of $h$. We substitute $\epsilon$ by $\frac{\epsilon}{2c\log n}$. Then, the approximation ratio is bounded as
@@
where $f$ is some computable function.
@@ C.2 Clique-Width
The following lemma corresponds to Lemma 7.5. The proof is the same as Lemma C.7.
Lemma C.12.
@@
Theorem 1.7 is also proved by the same proof as Theorem 1.2.

### v2

@@ 7 Clique-Width
In this section, we prove Theorems 6 and 7. Since the proofs are similar to that for Theorem 6, we provide most of the proof of Theorem 7 in Appendix C. Let $G$ be a vertex-colored graph with $n$ vertices. A VR-term $\mathcal{T}$ of $G$ represents a procedure that constructs $G$ by recursively applying the following operations [4]:
@@ 7.1 Definitions
for minimization problems. If $\mathsf{sat}(G,\varphi)=\emptyset$, we define $\mathsf{opt}_{w}[G,\varphi]=\mathtt{nil}$. For graphs $G$ and $\textsf{MSO}_{1}$ formulas $\varphi$ with $\mathsf{sat}(G,\varphi)\neq\emptyset$, our algorithm recursively computes the vertex set $\mathsf{DP}_{w}[G,\varphi]$ that approximately achieves the maximum in Equation (5) or minimum in Equation (6). When it is clear from the context, we omit the subscript $w$ and simply write $\mathsf{opt}[G,\varphi]$ and $\mathsf{DP}[G,\varphi]$. We perform the dynamic programming algorithm using the formulas defined in Theorem 31 over the parse tree of a term obtained from Lemma 30. Specifically, at every stage of the algorithm, it is guaranteed that each $X\in\mathsf{sat}(G,\varphi)$ satisfies $X\subseteq V(G)$. Furthermore, our algorithm is randomized. Thus, $\mathsf{DP}[G,\varphi]$ can be considered as a probability distribution over vertex sets in $\mathsf{sat}(G,\varphi)$.
@@ 7.4 Putting Together
Here we complete the proof of Theorem 6. Let $G=(V,E)$ be a graph with clique-width $\mathrm{cw}$. From Lemma 8, we can compute a VR-term $\mathcal{T}$ over a $k$-colored graph denoting $G$ with height $O(\log n)$. We perform the dynamic programming algorithm on the parse tree of $\mathcal{T}$. We remark here that operations other than disjoint union do not affect either the approximability or the Lipschitz continuity, by setting
@@
Proof.
The approximability bound is obtained by repeatedly applying Lemma 32 and $(1-\varepsilon)^{h}\geq 1-h\varepsilon$. The Lipschitzness bound is obtained by repeatedly applying Lemma 33. ∎
@@
Proof of Theorem 6.
The claim follows by substituting $\varepsilon$, $h$, and $k$ in Lemma 24 with $\frac{\varepsilon}{\Theta(\log n)}$, $O(\log n)$, and $O^{*}(2^{k})$, respectively. ∎
@@ Appendix A Omitted Proofs on Logic
It is easy to verify that the resulting parse tree represents $G$, and the height is at most $O(\log k+h)$. ∎
Proof of Corollary 14.
First we show the first statement of Corollary 14. Since $S^{\prime}\subseteq B\subseteq\mathsf{src}(G)$, $\mathsf{sat}(G,\psi\mathord{{\restriction}_{S\cup S^{\prime}}})$ and $\{S^{\prime}\}$ is separated. It is clear that families of vertex sets $(\mathsf{sat}(G,\psi\mathord{{\restriction}_{S\cup S^{\prime}}})\boxtimes\{S^{\prime}\})_{S^{\prime}\subseteq B}$ are disjoint.
Let $\psi(X)$ be a $\textsf{C}_{r}\textsf{MSO}^{q}$-formula obtained by applying the first statement of Theorem 12 to $\varphi(X)$ and $B$. Then, $\mathsf{fg}_{B}(G)\models\varphi(A)\iff G\models\psi(A)$ for any vertex set $A$. We show that $\mathsf{sat}(\mathsf{fg}_{B}(G),\varphi\mathord{{\restriction}_{S}})=\biguplus_{S^{\prime}\subseteq B}\mathsf{sat}(G,\psi\mathord{{\restriction}_{S\cup S^{\prime}}})\boxtimes\{S^{\prime}\}.$
@@
Next, we show the second statement of Corollary 14. From Theorem 12, there exists a family of tuples $\{\theta^{\prime}_{i}({X}),\psi^{\prime}_{i}({X})\}_{i\in[p^{\prime}]}$ of $\textsf{C}_{r}\textsf{MSO}^{q}$-formulas with a free variable ${X}$ such that, for any $k$-graphs $G$ and $H$,
@@ Appendix B Analysis of Softmax and Softmin
$\mathsf{TV}\left(\bm{c},\bm{c}^{\prime}\right)=\Pr\left[\bm{c}\in\left[\frac{2q}{x+\delta},\frac{2q}{x}\right]\right]=\frac{\frac{2q}{x}-\frac{2q}{x+\delta}}{\frac{2q}{x}-\frac{q}{x}}=2-\frac{2x}{x+\delta}=\frac{2\delta}{x+\delta}\leq\frac{2\delta}{x}.\qed$
@@
$=\left(1-\frac{\epsilon}{2}\right)\cdot\left(1-\frac{\epsilon}{2}\right)\max_{i\in[p]}x_{i}\geq(1-\epsilon)\max_{i\in[p]}x_{i}.\qed$
Proof of Lemma 17.
@@
$\mathsf{TV}\left(\mathsf{softargmax}^{\epsilon}_{i\in[p]}x_{i},\mathsf{softargmax}^{\epsilon}_{i\in[p]}x_{i}^{\prime}\right)\leq\frac{2\delta}{\max_{i\in[p]}x_{i}}+\frac{8\epsilon^{-1}\log(2p\epsilon^{-1})\delta}{\max_{i\in[p]}x_{i}}\leq\frac{10\epsilon^{-1}\log(2p\epsilon^{-1})\delta}{\max_{i\in[p]}x_{i}}.\qed$
@@
$\leq\left(\left(1+\frac{\epsilon}{2}\right)+\left(\frac{\epsilon}{2}\right)^{2}\cdot 2\right)\cdot\min_{i\in[p]}x_{i}\leq(1+\epsilon)\min_{i\in[p]}x_{i}.\qed$
Proof of Lemma 19.
@@
$\mathsf{TV}\left(\mathsf{softargmin}^{\epsilon}_{i\in[p]}x_{i},\mathsf{softargmin}^{\epsilon}_{i\in[p]}x_{i}^{\prime}\right)\leq\frac{2\delta}{\min_{i\in[p]}x_{i}}+\frac{8\epsilon^{-1}\log(2p\epsilon^{-1})\delta}{\min_{i\in[p]}x_{i}}\leq\frac{10\epsilon^{-1}\log(2p\epsilon^{-1})\delta}{\min_{i\in[p]}x_{i}}.\qed$
@@ C.1 Treewidth
Here, we provide the minimization version of the framework from Section 3. The algorithm is obtained simply by replacing $\mathsf{softargmax}$ with $\mathsf{softargmin}$. Some parts of the analysis are identical to that in Section 3. Specifically, for the introduce operation, exactly the same results can be obtained with exactly the same proof. Furthermore, by replacing $(1+\epsilon)$ with $(1-\epsilon)$ and reversing all inequalities in the proofs of Lemmas 20 and 22, we obtain the following.
Lemma 36.
@@
$\leq 30\alpha\epsilon^{-1}\log(2p\epsilon^{-1})\delta+\beta+\delta\leq 31\alpha\epsilon^{-1}\log(2p\epsilon^{-1})\delta+\beta.\qed$
The following lemma corresponds to Lemma 24, but slightly different.
Lemma 40.
@@
Proof.
The approximability bound is obtained by repeatedly applying Lemmas 36 and 37. The Lipschitzness bound is obtained by repeatedly applying Lemmas 38 and 39 together with $\alpha\leq(1+\epsilon)^{h}$. ∎
@@
Proof of Theorem 2.
We prove the theorem by substituting $\epsilon$, $h$, and $t$ in Lemma 40 with $\frac{\epsilon}{O(\log n)}$, $O(\log n)$, and $23t-$ 2, respectively. Let $c\in\mathbb{R}_{>0}$ be the number such that $c\log n$ is an upper bound of $h$. We substitute $\epsilon$ by $\frac{\epsilon}{2c\log n}$. Then, the approximation ratio is bounded as
@@
where $f$ is some computable function. ∎
@@ C.2 Clique-Width
The following lemma corresponds to Lemma 34. The proof is the same as Lemma 40.
Lemma 43.
@@
Theorem 7 is also proved by the same proof as Theorem 2.